OpenStudy (annie12m):

The price of a mobile phone decreased from $375 to $250 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 32% B. 33% C. 50% D. 83%

OpenStudy (annie12m):

@K12girlohva @lolacole12 @Liam2004

OpenStudy (k12girlohva):

This website can help you also freemathhelp.com/factoring-calculator.php

OpenStudy (k12girlohva):

http://www.freemathhelp.com/factoring-calculator.php

OpenStudy (lolacole12):

So decrease is 375-250=75 so convert 75/375 to percent percent means part out of 100 so x/100=x% divide 125/375=0.3333/1 multiply by 100/100=33/100=33% the answer is 33% or B

OpenStudy (k12girlohva):

Yes B is correct

OpenStudy (annie12m):

ty @lolacole12 @K12girlohva

OpenStudy (k12girlohva):

ANy1 DISAGBREE SPEAK NOW OR FOREVER HOLD UR PEACE

OpenStudy (anonymous):

well if you need help with this use this equation 375 - 250 is what percent of 375

OpenStudy (anonymous):

thus making b. correct

OpenStudy (anonymous):

I hope this helps you understand it more

OpenStudy (annie12m):

it did @Liam2004 ty

OpenStudy (lolacole12):

:/

OpenStudy (lolacole12):

gj

OpenStudy (anonymous):

see you just need to subtract for these types of questions and find what is the percentage